An elevator cab that weighs 27.8 kN moves upward. What is the tension in the cable if the cab’s speed is ___.

 

Find the mass of the elevator:

a.    At rest                                          

b.    Uniform motion

a.    Increasing at a rate of 1.2m/s2     

b.    Decreasing at the rate of 1.2m/s2

1
Expert's answer
2022-02-11T08:46:31-0500

Given:

"W=27.8*10^3\:\rm N"


The tension in cable

"T=W+ma"

a. "a=0"

"T=W=27.8\rm \: kN"

b. "a=0"

"T=W=27.8\rm \: kN"

c. "a=1.2\:\rm m/s^2"

"T=W+ma\\
=27.8\rm \: kN+\frac{27.8\rm \: kN}{9.8\:\rm N/kg}*1.2\:m/s^2=31200\:\rm N"

d. "a=-1.2\:\rm m/s^2"

"T=W+ma\\
=27.8\rm \: kN-\frac{27.8\rm \: kN}{9.8\:\rm N/kg}*1.2\:m/s^2=24396\:\rm N"